4. What volume of 10 M HCl and 3 M HCl should be mixed to get 1 litre of 6 M HCl solution?

 5. Magnesium carbonate reacts with sulphuric acid as shown below:

MgCO3 + H2SO4 →MgSO4 + H2O + CO2

i. What weight of sulphuric acid will be required to dissolve three grams of

ii. Magnesium carbonate?

1
Expert's answer
2020-07-22T05:41:52-0400

Question 4

lets represent the molarity of the acid using letters, i.e. for 10M=M1 and 3M=M2

molarity of a mixture is determined from the formula:

(M1V1 + M2V2) / (V1 + V2) = M3 

hence, replacing the values given, we obtain the result below:

[10x + 3(1-x)] / 1 = 6 (we then proceed by solution the unknown)

10x - 3x + 3 = 6 

7x = 3 

x = 3 / 7 = 0.43

Therefore, for 10M HCl = 0.43L

for 3 M HCl = 1-0.43 = 0.57 L

Question 5

MgCO3 + H2SO4 →MgSO4 + H2O + CO2

Mass of magnesium carbonate = 3g

using reaction mole ratios, we can determine equivalent moles of sulfuric acid.

The formula for calculating moles = mass/molar mass of a salt

molar mass of magnesium carbonate is derived from addition of molar mass of constituent atoms present

molar mass of magnesium carbonate = 24+12+48 = 84

moles of the salt = 3/84 = 0.0357 moles

reaction mole ratio of the salt and sulfuric acid 1:1

hence, the number of moles of sulfuric acid = 0.0357 moles

from the formula for calculating moles, we can determine the mass of the acid

therefore, mass of the acid = 0.0357*98 = 3.4986g